The sum of the coordinates of the solution to the system
2x-y=7
3x-2y=11
is?

Answer by Alan3354(21605)


The sum of the coordinates of the solution to the system
2x-y=7
3x-2y=11
is?
------------
What can that mean? Sum of coordinates?
So solve for x and y by substition:
From eqn 1, y = 2x-7
Sub into eqn 2:
3x - 2*(2x-7) = 11
3x -4x +14 = 11
-x = -3
x = 3
Sub into either eqn
2*(3) - y = 7
6 -y = 7
y = -1